OpenStudy (anonymous):

Mulitply the polynomials [(x+y)-8]^2

OpenStudy (anonymous):

you remember \[(a-b)^2=a^2-2ab+b^2\] so now [(x+y)-8]^2=(x+y)^2-2(x+y)(8)+64) =x^2+2xy+y^2-16(x+y)+64 =x^2+2xy+y^2-16x-16y+64